OpenStudy (anonymous):

When a net force of 17.0 newtons is applied to a dictionary placed on a frictionless table, it accelerates by 3.75 meters/secondÆ. What is the mass of the book? A. 4.00 B. 4.53 C. 5.00 D. 6.00 IM TRYING TO FINISH THIS SO I CAN GRADUATE I NEED TO HURRY

OpenStudy (anonymous):

I HAVE TO GET 4 RIGHT TO PASS

OpenStudy (stormfire1):

\[a= 3.75m/s^2\]\[F = 17.0N\]\[F=ma\]\[17.0N=m(3.75 \frac{m}{s^2})\]\[\frac{17.0N}{3.75 \frac{m}{s^2}}=a=?\]

OpenStudy (anonymous):

THE ANSWER IS A?

OpenStudy (anonymous):

idk the answer

OpenStudy (stormfire1):

what is 17 divided by 3.75?

OpenStudy (anonymous):

4.53

OpenStudy (stormfire1):

bingo.

OpenStudy (anonymous):

4.53

OpenStudy (anonymous):

f=ma f=17 a=3.75 m=? f=ma m=f/a m=17/3.75